# Why Policy Gradient Algorithms Work for Undiscounted Total-Reward MDPs

### v1

@@ 5 Convergence of projected policy gradient
$\Pi_{\alpha}=\{\pi\,|\,\pi(s\,|\,a)\geq\alpha\text{ for all }s\in\mathcal{S},\,a\in\mathcal{A}\}$
with $\alpha\in(0,1)$. For evaluating $\nabla V_{\mu}^{\pi_{k}}$, Theorem 1 applied to the direct parametrization setp yields $\nabla V_{\mu}^{\pi_{k}}(s,a)=\delta_{\mu}^{\pi_{\theta}}(s)Q^{\pi_{\theta}}(s,a)$ for $\pi\in\Pi_{\alpha}$.
@@
Lemma 6 (Smoothness of value function).
Under Assumption 1, for $\pi,\pi^{\prime}\in\Pi_{\alpha}$,
@@
Define $V^{\pi^{\star}_{\alpha}}=\max_{\pi\in\Pi_{\alpha}}V^{\pi}$ and $V^{\pi^{\star}_{\alpha}}_{\mu}=\mathbb{E}_{s\sim\mu}[V^{\pi^{\star}_{\alpha}}(s)]$. Using the Lemma 6, we obtain the following convergence result of the projected policy gradient algorithm.
Theorem 2.
Under Assumption 1, for $\alpha\in(0,1)$, $\pi_{0}\in\Pi_{\alpha}$, and given $\mu$ with full support, the projected policy gradient algorithm with step size $\eta=\dfrac{1}{2RC^{2}_{\alpha}(C_{\alpha}+1)|\mathcal{A}|}$ generates a sequence of policies $\{\pi_{k}\}^{\infty}_{k=1}$ satisfying
@@
We defer the proofs to Appendix B, but we quickly note that the proof strategy closely follows Xiao (2022), which considers the discounted reward setup and uses the (non-transient) visitation measure.
Theorem 1 shows that $V_{\mu}^{\pi_{k}}\rightarrow V_{\mu}^{\pi^{\star}_{\alpha}}$ with a sublinear rate, and since $V^{\pi^{\star}_{\alpha}}_{\mu}\rightarrow V^{\star}_{+,\mu}$ as $\alpha\rightarrow 0$, we can define an iteration complexity for finding an $\epsilon$-optimal policy by projecting onto $\Pi_{\alpha}$ with the value of $\alpha$ chosen to be a function of $\epsilon$ with nonnegative reward.
Corollary 1.
Assume the rewards are nonnegative. For any given $\epsilon\in(0,1)$ and $\mu$ with full support, set $\alpha=\frac{\epsilon}{2|\mathcal{S}||\mathcal{A}|\left\|{\delta^{\pi^{\star}}_{\mu}}\right\|_{\infty}\left\|{Q^{\pi^{\star}}}\right\|_{\infty}}$ and let the step size be $\eta=\dfrac{1}{2RC^{2}_{\alpha}(C_{\alpha}+1)|\mathcal{A}|}$. Then, under Assumption 1, for $\pi_{0}\in\Pi_{\alpha}$, the iterates of projected policy gradient $\pi_{k}$ are $\epsilon$-optimal policy for
@@
In Theorem 2, we establish convergence to $\epsilon$-optimality with $\epsilon=V_{\mu}^{\star}-V_{\mu}^{\pi^{\star}_{\alpha}}$, which can be made arbitrarily small by taking $\alpha>0$ to be small. However, we do not have convergence to exact optimality as $k\to\infty$, where $k$ is the iteration count. Moreover, the convergence rate is sublinear, and the constant factor depends on the sizes of the state and action spaces, which may be quite large. These shortcomings are addressed by the analysis of the natural policy gradient method presented in the next section.
@@ 6 Convergence of natural policy gradient
$\pi_{\theta}(a\,|\,s)=\frac{\exp(\theta_{s,a})}{\sum_{a^{\prime}}\exp(\theta_{s,a}).}$
@@
$\begin{aligned} F_{\mu}(\theta^{k})&=\sum_{s\in\mathcal{S}}\delta^{\pi_{\theta}}_{\mu}(s)\mathbb{E}_{a\sim\pi_{\theta}(\cdot\mid s)}\Big[\nabla_{\theta}\log\pi_{\theta}(a\mid s)\,\big(\nabla_{\theta}\log\pi_{\theta}(a\mid s)\big)^{\top}\Big]\\ \theta^{k+1}&=\theta^{k}+\eta_{k}F_{\mu}(\theta^{k}\big)^{\dagger}\,\nabla_{\theta}V^{\pi_{k}}_{\mu}\end{aligned}\qquad\text{for }k=0,1,\dots,$
where $\dagger$ denotes the Moore–Penrose pseudoinverse. It is known that natural policy gradient algorithm can also be expressed as
$\pi_{k+1}(a\,|\,s)=\pi_{k}(a\,|\,s)\frac{\exp(\eta_{k}Q^{\pi_{k}}(s,a))}{z_{s}^{k}}\propto\pi_{0}(a\,|\,s)\exp\Big(\sum_{i=0}^{k}\eta_{i}Q^{\pi_{n}}(s,a)\Big)$
@@
is a normalization factor. In the online learning literature, this update rule is also known as multiplicative weights updates (Freund & Schapire, 1997) and the update rule does not depend on the initial state distribution $\mu$, as the pseudoinverse of the Fisher information removes this dependency.
@@ 6.1 Sublinear convergence with constant step size
Lemma 7.
Under Assumption 1, for $\pi_{0}\in\Pi_{+}$ and given $\mu$ with full support, the natural policy gradient with constant step size $\eta>0$ generates a sequence of policies $\{\pi_{k}\}^{\infty}_{k=1}$ satisfying
@@
where $\mathrm{KL}(p,q)=\sum^{n}_{i=1}p_{i}\log(p_{i}/q_{i})$ for $p,q\in\mathcal{M}(\mathcal{A})$, and also define $\left\|{(I-T^{\pi})^{-1}}\right\|_{\infty}=C_{\pi}<\infty.$ Combining Lemmas 5 and 7, we obtain the following sublinear convergence.
Theorem 3.
Under Assumption 1, for $\pi_{0}\in\Pi_{+}$ and given $\mu$ with full support, the natural policy gradient with constant step size $\eta>0$ generates a sequence of policies $\{\pi_{k}\}^{\infty}_{k=1}$ satisfying
$V_{\mu}^{\pi}-V_{\mu}^{\pi_{k}}\leq\frac{1}{k+1}\left(\frac{\mathrm{KL}_{\delta_{\mu}^{\pi}}(\pi,\pi_{0})}{\eta}+C_{\pi}(\left\|{V^{\star}_{+}}\right\|_{\infty}+\left\|{V^{\pi_{0}}}\right\|_{\infty})\right).$
@@
Appendix C provides the proof, which closely follows Xiao (2022).
@@
Corollary 2.
Assume the rewards are nonnegative. Under Assumption 1, for $\pi_{0}\in\Pi_{+}$ and given $\mu$ with full support, the natural policy gradient algorithm with constant step size $\eta>0$ generates a sequence of policies $\{\pi_{k}\}^{\infty}_{k=1}$ satisfying
@@
Next, we present the linear convergence rate of the natural policy gradient algorithm with adaptive step size. For a given $\mu$ with full support, define
$\vartheta^{\pi}_{\mu}=\left\|\frac{\delta_{\mu}^{\pi}}{\mu}\right\|_{\infty}\in[1,\infty),$
which represent the distribution mismatch between $\mu$ and $\delta_{\mu}^{\pi}$.
Theorem 4.
Under Assumption 1, for $\pi_{0}\in\Pi_{+}$ and $\mu$ with full support, the natural policy gradient algorithm with step sizes $(\vartheta^{\pi}_{\mu}-1)\eta_{k+1}\geq\vartheta^{\pi}_{\mu}\eta_{k}>0$ generates a sequence of policies $\{\pi_{k}\}^{\infty}_{k=1}$ satisfying
@@
Again, the convergence rate is independent of the size of the state or action space, and we can strengthen the convergence result if we assume the rewards are nonnegative.
Corollary 3.
Assume the rewards are nonnegative. Under Assumption 1, for $\pi_{0}\in\Pi_{+}$ and given $\mu$ with full support, the natural policy gradient algorithm with step size $(\vartheta^{\pi}_{\mu}-1)\eta_{k+1}\geq\vartheta^{\pi}_{\mu}\eta_{k}>0$ generates a sequence of policies $\{\pi_{k}\}^{\infty}_{k=1}$ satisfying
$V_{\mu}^{\star}-V_{\mu}^{\pi_{k}}\;\leq\;\left(1-\frac{1}{\vartheta^{\pi^{\star}}_{\mu}}\right)^{\!k}\left(V_{\mu}^{\star}-V_{\mu}^{\pi_{0}}+\frac{\mathrm{KL}_{\delta_{\mu}^{\pi^{\star}}}(\pi^{\star},\pi_{0})}{\eta_{0}(\vartheta^{\pi^{\star}}_{\mu}-1)}\right).$

### v2

@@ 5 Convergence of projected policy gradient
$\Pi_{\alpha}=\{\pi\,|\,\pi(a\,|\,s)\geq\alpha\text{ for all }s\in\mathcal{S},\,a\in\mathcal{A}\}$
with $\alpha\in(0,1/|\mathcal{A}|)$. For evaluating $\nabla V_{\mu}^{\pi_{k}}$, Theorem 4.7 applied to the direct parametrization setup yields $\nabla V_{\mu}^{\pi_{k}}(s,a)=\delta_{\mu}^{\pi_{\theta}}(s)Q^{\pi_{\theta}}(s,a)$ for $\pi\in\Pi_{\alpha}$.
@@
Lemma 5.1 (Smoothness of value function).
Under Assumption 2.1, for $\pi,\pi^{\prime}\in\Pi_{\alpha}$,
@@
Define $V^{\pi^{\star}_{\alpha}}=\max_{\pi\in\Pi_{\alpha}}V^{\pi}$ and $V^{\pi^{\star}_{\alpha}}_{\mu}=\mathbb{E}_{s\sim\mu}[V^{\pi^{\star}_{\alpha}}(s)]$. Using the Lemma 5.1, we obtain the following convergence result of the projected policy gradient algorithm.
Theorem 5.2.
Under Assumption 2.1, for $\alpha\in(0,1)$, $\pi_{0}\in\Pi_{\alpha}$, and given $\mu$ with full support, the projected policy gradient algorithm with step size $\eta=\dfrac{1}{2RC^{2}_{\alpha}(C_{\alpha}+1)|\mathcal{A}|}$ generates a sequence of policies $\{\pi_{k}\}^{\infty}_{k=1}$ satisfying
@@
We defer the proofs to Appendix B, but we quickly note that the proof strategy closely follows (Xiao, 2022), which considers the discounted reward setup and uses the (non-transient) visitation measure.
Theorem 5.2 shows that $V_{\mu}^{\pi_{k}}\rightarrow V_{\mu}^{\pi^{\star}_{\alpha}}$ with a sublinear rate, and since $V^{\pi^{\star}_{\alpha}}_{\mu}\rightarrow V^{\star}_{+,\mu}$ as $\alpha\rightarrow 0$, if we choose a sufficiently small $\alpha$ such that $V_{\mu,+}^{\star}-V_{\mu}^{\pi_{\alpha}}<\frac{\epsilon}{2}$, we can obtain policy $\pi$ satisfying $V_{\mu,+}^{\star}-V_{\mu}^{\pi}<\epsilon$. Moreover, we can define an iteration complexity for finding an $\epsilon$-optimal policy by projecting onto $\Pi_{\alpha}$ with the value of $\alpha$ chosen to be a function of $\epsilon$ with nonnegative reward.
Corollary 5.3.
Assume the rewards are nonnegative. For any given $\epsilon\in(0,1)$ and $\mu$ with full support, set $\alpha=\frac{\epsilon}{2|\mathcal{S}||\mathcal{A}|\left\|{\delta^{\pi^{\star}}_{\mu}}\right\|_{\infty}\left\|{Q^{\pi^{\star}}}\right\|_{\infty}}$ and let the step size be $\eta=\dfrac{1}{2RC^{2}_{\alpha}(C_{\alpha}+1)|\mathcal{A}|}$. Then, under Assumption 2.1, for $\pi_{0}\in\Pi_{\alpha}$, the iterates of projected policy gradient $\pi_{k}$ are $\epsilon$-optimal policy for
@@
In Theorem 5.2, we establish convergence to $\epsilon$-optimality with $\epsilon=V_{\mu}^{\star}-V_{\mu}^{\pi^{\star}_{\alpha}}$, which can be made arbitrarily small by taking $\alpha>0$ to be small. However, we do not have convergence to exact optimality as $k\to\infty$, where $k$ is the iteration count. Moreover, the convergence rate is sublinear, and the constant factor depends on the sizes of the state and action spaces, which may be quite large. These shortcomings are addressed by the analysis of the natural policy gradient method presented in the next section.
@@ 6 Convergence of natural policy gradient
$\pi_{\theta}(a\,|\,s)=\frac{\exp(\theta_{s,a})}{\sum_{a^{\prime}}\exp(\theta_{s,a^{\prime}}).}$
@@
$\begin{aligned} \theta^{k+1}&=\theta^{k}+\eta_{k}F_{\mu}(\theta^{k}\big)^{\dagger}\,\nabla_{\theta}V^{\pi_{k}}_{\mu}\end{aligned}\quad\text{for }k=0,1,\dots,$
where $F_{\mu}(\theta^{k})=\sum_{s\in\mathcal{S}}\delta^{\pi_{\theta}}_{\mu}(s)\mathbb{E}_{a\sim\pi_{\theta}(\cdot\mid s)}\Big[\nabla_{\theta}\log\pi_{\theta}(a\mid s)\,\big(\nabla_{\theta}\log\pi_{\theta}(a\mid s)\big)^{\top}\Big]$ and $\dagger$ denotes the Moore–Penrose pseudoinverse. It is known that natural policy gradient algorithm can also be expressed as
$\pi_{k+1}(a\,|\,s)$ $=\pi_{k}(a\,|\,s)\frac{\exp(\eta_{k}Q^{\pi_{k}}(s,a))}{z_{s}^{k}}$
$\propto\pi_{0}(a\,|\,s)\exp\Big(\sum_{i=0}^{k}\eta_{i}Q^{\pi_{i}}(s,a)\Big)$
@@
is a normalization factor. In the online learning literature, this update rule is also known as multiplicative weights updates (Freund and Schapire, 1997) and the update rule does not depend on the initial state distribution $\mu$, as the pseudoinverse of the Fisher information removes this dependency.
@@ 6.1 Sublinear convergence with constant step size
Lemma 6.1.
Under Assumption 2.1, for $\pi_{0}\in\Pi_{+}$ and given $\mu$ with full support, the natural policy gradient with constant step size $\eta>0$ generates a sequence of policies $\{\pi_{k}\}^{\infty}_{k=1}$ satisfying
@@
where $\mathrm{KL}(p,q)=\sum^{n}_{i=1}p_{i}\log(p_{i}/q_{i})$ for $p,q\in\mathcal{M}(\mathcal{A})$, and also define $\left\|{(I-T^{\pi})^{-1}}\right\|_{\infty}=C_{\pi}<\infty.$ Combining Lemmas 4.6 and 6.1, we obtain the following sublinear convergence.
Theorem 6.2.
Under Assumption 2.1, for any $\pi,\pi_{0}\in\Pi_{+}$ and given $\mu$ with full support, the natural policy gradient with constant step size $\eta>0$ generates a sequence of policies $\{\pi_{k}\}^{\infty}_{k=1}$ satisfying
$V_{\mu}^{\pi}-V_{\mu}^{\pi_{k}}\leq\frac{1}{k+1}\left(\frac{\mathrm{KL}_{\delta_{\mu}^{\pi}}(\pi,\pi_{0})}{\eta}+2C_{\pi}\left\|{V^{\star}_{+}}\right\|_{\infty}\right).$
@@
Appendix C provides the proof, which closely follows (Xiao, 2022).
@@
Corollary 6.3.
Assume the rewards are nonnegative. Under Assumption 2.1, for any $\pi,\pi_{0}\in\Pi_{+}$ and given $\mu$ with full support, the natural policy gradient algorithm with constant step size $\eta>0$ generates a sequence of policies $\{\pi_{k}\}^{\infty}_{k=1}$ satisfying
@@
Next, we present the linear convergence rate of the natural policy gradient algorithm with adaptive step size. For a given $\mu$ with full support, define $\vartheta^{\pi}_{\mu}=\left\|\frac{\delta_{\mu}^{\pi}}{\mu}\right\|_{\infty}\in[1,\infty),$ which represent the distribution mismatch between $\mu$ and $\delta_{\mu}^{\pi}$.
Theorem 6.4.
Under Assumption 2.1, for $\pi_{0}\in\Pi_{+}$ and $\mu$ with full support, the natural policy gradient algorithm with step sizes $(\vartheta^{\pi}_{\mu}-1)\eta_{k+1}\geq\vartheta^{\pi}_{\mu}\eta_{k}>0$ generates a sequence of policies $\{\pi_{k}\}^{\infty}_{k=1}$ satisfying
@@
As the distribution mismatch decreases, we can see that the natural policy gradient converges faster. Again, the convergence rate is independent of the size of the state or action space, and we can strengthen the convergence result if we assume the rewards are nonnegative.
Corollary 6.5.
Assume the rewards are nonnegative. Under Assumption 2.1, for $\pi_{0}\in\Pi_{+}$ and given $\mu$ with full support, the natural policy gradient algorithm with step size $(\vartheta^{\pi^{\star}}_{\mu}-1)\eta_{k+1}\geq\vartheta^{\pi^{\star}}_{\mu}\eta_{k}>0$ generates a sequence of policies $\{\pi_{k}\}^{\infty}_{k=1}$ satisfying
$V_{\mu}^{\star}-V_{\mu}^{\pi_{k}}\leq\left(1-\frac{1}{\vartheta^{\pi^{\star}}_{\mu}}\right)^{\!k}\left(V_{\mu}^{\star}-V_{\mu}^{\pi_{0}}+\frac{\mathrm{KL}_{\delta_{\mu}^{\pi^{\star}}}(\pi^{\star},\pi_{0})}{\eta_{0}(\vartheta^{\pi^{\star}}_{\mu}-1)}\right)$
@@
7 Stochastic natural policy gradient
In this section, we extend the analysis of the natural policy gradient to the sampling setting in which the transition probabilities are unknown. Specifically, we assume access to a generative model (Kearns and Singh, 1998), which provides independent samples of the next state for any given state and action.
7.1 Approximating $Q$-function with a generative model
With generative model, for a given policy $\pi$ and any state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$, we can generate independent trajectories of horizon $H$, i.e.,
$\left\{\bigl(s_{0},a_{0}\bigr),\bigl(s_{1},a_{1}\bigr),\ldots,\bigl(s_{H-1},a_{H-1}\bigr)\ \middle|\ s_{0}=s,\ a_{0}=a\right\}.$
With this samples, we consider following $Q$-estimator to approximate state-action value function.
$\tilde{Q}^{\pi}=\frac{1}{N}\sum^{N}_{j=1}\tilde{Q}_{j}^{\pi}\quad\text{where}\quad\tilde{Q}_{j}^{\pi}(s_{0},a_{0})=\sum^{H}_{i=0}r_{j}(s_{i},a_{i})$
for all $s_{0}\in\mathcal{S}$, $a_{0}\in\mathcal{A}$.
Next, we define transient half-life for given $\pi$ as
$t_{\frac{1}{2},\pi}=\operatorname{argmin}_{t\geq 1}\left\|{(T^{\pi})^{t}}\right\|_{\infty}\leq\frac{1}{2},$
where $\left\|{\cdot}\right\|_{\infty}$ denotes the maximum absolute row sum of the matrix. This transient half-life is always finite due to Fact 1.
We now present sample complexity on the $Q$-estimator with generative model.
Theorem 7.1.
Let $\epsilon>0$ and $\delta>0$. Under Assumption 2.1, for given $\pi\in\Pi$, with $1-\delta$ probability, $\left\|{\tilde{Q}^{\pi}-Q^{\pi}}\right\|_{\infty}\leq\epsilon$ with sample complexity
$\widetilde{\mathcal{O}}\left(\frac{t_{\frac{1}{2},\pi}^{3}R^{2}|\mathcal{S}||\mathcal{A}|}{\epsilon^{2}}\right)$
where $\widetilde{\mathcal{O}}$ ignores all logarithmic factors.
Note that this holds for arbitrary $\pi\in\Pi$, and if we priori know classification of states, $|\mathcal{S}|$ in sample complexity of Theorem 7.1 could be reduced to number of transient states.
7.2 Sample complexity of stochastic natural policy gradient
By approximating $Q$-function, stochastic natural policy gradient can be expressed as
$\tilde{Q}^{\pi_{k}}(s_{0},a_{0})=\frac{1}{N}\sum^{N}_{j=1}\sum^{H}_{i=0}r_{j}(a_{i},s_{i})$
$\pi_{k+1}(a\,|\,s)=\pi_{k}(a\,|\,s)\frac{\exp(\eta_{k}\tilde{Q}^{\pi_{k}}(s,a))}{z_{s}^{k}}$
for all $s_{0},s\in\mathcal{S}$, $a,a_{0}\in\mathcal{A}$, and $k=0,1,\dots$, where $z^{k}_{s}=\sum_{a\in\mathcal{A}}\pi_{k}(a\,|\,s)\exp(\eta_{k}Q^{\pi_{k}}(s,a))$. Note that unlike projected policy gradient, natural policy gradient only requires information of state-action value function.
@@ 7.2 Sample complexity of stochastic natural policy gradient
Now we present sample complexity of stochastic natural policy gradient. We defer the proofs to Appendix D, but we note that the proof strategy closely follows (Xiao, 2022).
Theorem 7.2.
Let $\epsilon>0$ and $\delta>0$. Under Assumption 2.1, for any $\pi,\pi_{0}\in\Pi_{+}$ and given $\mu$ with full support, with probability $1-\delta$, the iterate of stochastic natural policy gradient with constant step size $\eta>0$ and $K=2\Big(\frac{\text{KL}_{\rho_{\mu}^{\pi}}(\pi,\pi_{0})}{\eta}+2C_{\pi}\left\|{V^{\star}_{+}}\right\|_{\infty}\Big)/\epsilon$ satisfies $V_{\mu}^{\pi}-V_{\mu}^{\pi_{K}}\leq\epsilon$ with sample complexity
$\widetilde{\mathcal{O}}\left(\frac{t_{\frac{1}{2}}^{3}(\vartheta^{\pi}_{\mu})^{2}R^{2}C_{\pi}^{5}\left(\frac{\text{KL}_{\delta_{\mu}^{\pi}}(\pi,\pi_{0})}{\eta}\right)^{3}\left\|{V^{\star}_{+}}\right\|_{\infty}^{3}|\mathcal{S}||\mathcal{A}|}{\epsilon^{5}}\right)$
and with adaptive step size $(\vartheta^{\pi}_{\mu}-1)\eta_{k+1}\geq\vartheta^{\pi}_{\mu}\eta_{k}>0$ and $K=\log\left(2\left(V_{\mu}^{\pi}-V_{\mu}^{\pi_{0}}+\frac{\text{KL}_{\delta^{\pi}_{\mu}}(\pi,\pi_{0})}{\eta_{0}(\vartheta^{\pi}_{\mu}-1)}\right)/\epsilon\right)$ satisfies $V_{\mu}^{\pi^{\star}}-V_{\mu}^{\pi_{K}}\leq\epsilon$ with sample complexity
$\widetilde{\mathcal{O}}\left(\frac{t_{\frac{1}{2}}^{3}(\vartheta^{\pi}_{\mu})^{2}C_{K}^{2}R^{2}|\mathcal{S}||\mathcal{A}|}{\epsilon^{2}}\right)$
where $\widetilde{\mathcal{O}}$ ignores all logarithmic factors, $C_{K}=\max_{0\leq k\leq K-1}\{C_{\pi_{i}},C_{\pi}\}$, and $t_{\frac{1}{2}}=\max_{0\leq k\leq K-1}t_{\frac{1}{2},\pi_{k}}$.
Note that sample complexity of adaptive step size $\widetilde{\mathcal{O}}(1/\epsilon^{2})$ is more efficient that of constant step size $\widetilde{\mathcal{O}}(1/\epsilon^{5})$. We can also derive similar rates when the assume the rewards are nonnegative as follows.
Corollary 7.3.
Assume the rewards are nonnegative. Let $\epsilon>0$ and $\delta>0$. Under Assumption 2.1, for $\pi_{0}\in\Pi_{+}$ and given $\mu$ with full support, with probability $1-\delta$, the policy error of stochastic natural policy gradient with constant step size $\eta>0$ and $K=2\Big(\frac{\text{KL}_{\rho_{\mu}^{\pi^{\star}}}(\pi^{\star},\pi_{0})}{\eta}+C_{\pi^{\star}}\left\|{V^{\star}}\right\|_{\infty}\Big)/\epsilon$ satisfies $V_{\mu}^{\star}-V_{\mu}^{\pi_{K}}\leq\epsilon$ with sample complexity
$\widetilde{\mathcal{O}}\left(\frac{t_{\frac{1}{2}}^{3}(\vartheta^{\pi^{\star}}_{\mu})^{2}R^{2}C_{\pi^{\star}}^{5}\left(\frac{\text{KL}_{\delta_{\mu}^{\pi^{\star}}}(\pi^{\star},\pi_{0})}{\eta}\right)^{3}\left\|{V^{\star}}\right\|_{\infty}^{3}|\mathcal{S}||\mathcal{A}|}{\epsilon^{5}}\right)$
and with adaptive step size $(\vartheta^{\pi^{\star}}_{\mu}-1)\eta_{k+1}\geq\vartheta^{\pi^{\star}}_{\mu}\eta_{k}>0$ and $K=\log\left(2\left(V_{\mu}^{\star}-V_{\mu}^{\pi_{0}}+\frac{\text{KL}_{\delta^{\pi^{\star}}_{\mu}}({\pi^{\star}},\pi_{0})}{\eta_{0}(\vartheta^{\pi}_{\mu}-1)}\right)/\epsilon\right)$ satisfies $V_{\mu}^{\pi^{\star}}-V_{\mu}^{\pi_{K}}\leq\epsilon$ with sample complexity
$\widetilde{\mathcal{O}}\left(\frac{t_{\frac{1}{2}}^{3}(\vartheta^{\pi^{\star}}_{\mu})^{2}C_{K}^{2}R^{2}|\mathcal{S}||\mathcal{A}|}{\epsilon^{2}}\right)$
for any optimal policy $\pi^{\star}$ where $\widetilde{\mathcal{O}}$ ignores all logarithmic factors, $C_{K}=\max_{0\leq k\leq K-1}\{C_{\pi_{i}},C_{\pi^{\star}}\}$, and $t_{\frac{1}{2}}=\max_{0\leq k\leq K-1}t_{\frac{1}{2},\pi_{k}}$.